# Capture Stage Environments: A Guide to Better Matting

### v1

@@ Abstract
Capture stages are high-end sources of state-of-the-art recordings for downstream applications in movies, games, and other media. One crucial step in almost all pipelines is the matting of images to isolate the captured performances from the background. While common matting algorithms deliver remarkable performance in other applications like teleconferencing and mobile entertainment, we found that they struggle significantly with the peculiarities of capture stage content. The goal of our work is to share insights into those challenges as a curated list of those characteristics along with a constructive discussion for proactive intervention and present a guideline to practitioners for an improved workflow to mitigate unresolved challenges. To this end, we also demonstrate an efficient pipeline to adapt state-of-the-art approaches to such custom setups without the need of extensive annotations, both offline and real-time. For an objective evaluation, we propose a validation methodology based on a leading diffusion model that highlights the benefits of our approach.
@@ 1 Introduction
Matting is a fundamental technique in the field of computer vision, playing an important role in tasks requiring precise foreground-background separation. Mathematically, matting can be expressed by a linear combination of fore- and background images $F$ and $B$, enabling an alpha blending between foreground and background,
@@
Over the last decade, deep learning methods have driven the development of state-of-the-art matting techniques that are capable of fine-grained transparency estimation, which enabled the application of matting techniques in a wide range of fields, including augmented reality, film production. A particularly important application domain lies within capture stages that provide multi-camera setups for recording high-quality data in controlled environments. A robust matting approach is fundamental to reconstructing high-fidelity content as manual interaction is impractical due to the extensive volume of data involved. However, achieving such a matting in these setups remains a significant challenge, as state-of-the-art methods [5, 28, 24] typically rely on predefined trimaps or require large-scale labeled datasets for training, both of which are impractical for capture stage setups: While trimaps require unique and highly-accurate generation for every frame of every camera, existing matting datasets are limited in their generalizability to such unique environments, leading to incorrect matting results. Fortunately, it is the characteristics of capture stages, such as static backgrounds and controlled lighting, that enable the application of the empty stage as a reference for matting, providing additional information to support the matting process. However, this approach introduces new challenges, as reflective materials (e.g., structures within the stage) and soft shadows cast by the captured subject can significantly alter the appearance of the stage in each shot, complicating the matting workflow, especially given the persistent lack of high-quality training data.
In this work, we address these challenges by discussing the potential failure cases in capture stage setups, providing a guideline for improved matting in these environments, and improving deep matting techniques through a pipeline specifically designed for such capture stage setups. In particular, we place a special focus on real-time matting, which is highly relevant for streaming and telepresence applications. In this context, we discuss a two-phase pipeline based on a student-teacher concept. We modified an offline matting model to apply background matting and adapted it to process capture stage data using coarse user-drawn annotations in the first phase. This adapted model serves as a teacher in the second phase, guiding a real-time matting model to generate high-quality mattes for capture stage data by learning from the refined outputs of the teacher. We evaluate both models using a hand-labeled trimap-based diffusion model on a hold-out validation set. In summary, our contributions are:
@@
  • We validate our approach by comparing its accuracy against a precise, trimap-based matting method and present the effects of accurate matting in a NeRF-based downstream application, showing improvements in rendering quality (Sec. 5)
@@ 2 Related Work
With the rise of deep learning, Xu et al. [27] introduced a deep image matting model and released a dataset enabling training of models for alpha mask prediction. Since then, numerous works applied learning methods for deep image matting, many of which rely on pre-defined trimaps [28, 5], that often present a bottleneck due to the manual labelling required for their creation. Zhang et al. [31] addressed this challenge by learning trimaps from sparse user input, while others [17] developed a student-teacher model, letting a trimap-free student model learn from a trimap-guided teacher. Hachmann et al. [4] replaced the trimap completely by predicting the matting based on successive images of a scene with changing background colors. To facilitate the matting task, some works built upon segmentation models [11, 29], or used them for additional annotation [8], while another line of research used attention mechanisms and transformer models for matting [28, 2], e.g. by using cross-attention to capture object contours [26]. As diffusion models gained popularity, they have been increasingly utilized in the field of deep image matting [25, 3].
@@
Several works proposed real-time approaches for image and video matting to support certain down-stream applications like live-streaming. In this context, Lin et al. [14] presented a real-time matting model that uses additional background information for image matting, while other methods [7, 21, 15] solely rely on the image containing the foreground object. Among these methods, Ke et al. [7] proposed ModNet, a real-time matting model specifically designed for portrait matting, addressing flickering issues caused by the lack of temporal coherence in video matting. BiMatting [21] applies mask binarization to meet real-time requirements, while a separate recurrent-network solution [15] enhances robustness for continuous video matting. In this context, Huynh et al. [6] proposed a mask-guided method for human instance matting, while Lin et al. [13] developed a transformer-based approach for matting humans in unstructured video environments. For a comprehensive overview of deep image matting methods, please refer to the survey by Li et al. [10].
@@ 3 Problem Setup and Challenges
$I_{i}=\alpha_{i}F_{i}+(1-\alpha_{i})\,\bigl{(}B_{i}+\delta(F)\bigr{)}+\varepsilon_{i},\quad\alpha_{i}\in[0,1],$ (2)
whereby $\delta(F)$ represents alterations to $I_{i}$ caused by the influence of the foreground subjects shadows, reflections, or color bleed, while $\varepsilon$ represents sensor noise and other unmodeled artifacts.
Even in highly controlled studio conditions, minor deviations between the “ideal” background $B$ and the actual recordings of the stage with a foreground subject present are inevitable. These background alterations in $I$ are reinforced in particular by reflective materials such as the housings of cameras, lenses or metal frames, which reflect light and the subject. Shadows cast by the subject further alter the appearance of the background, whereby soft shadows occur especially at the feet of a subject, as shown in Fig. 2. Reflections and shadows make the pre-captured background information unreliable, leading to ambiguity in distinguishing the fore- and background. These deviations in the background can substantially affect the accuracy of alpha estimation, as the matting algorithm may fail to attribute background changes correctly to either the foreground layer $F$ or the background layer $B+\delta(F)$.
Like all camera setups, capture stages are subject to noise $\varepsilon$ which effectively alters the appearance of the background in each recording. Noise degrades the edges of objects, making pixel-accurate distinctions difficult, misleading a matting network into making incorrect predictions. While longer exposure times reduce noise, they also increase undesirable motion blur, particularly when capturing a moving person. In a capture stage setup, this can either require uncomfortably bright lights or high gain levels which further increase the noise and create more pronounced differences in the background.
@@
Figure 3: Ambiguities in foreground and background information can be caused by lighting in the background (left) and dark-on-dark areas (right).
@@ 4.1 Start with a Better Setup
The perhaps most crucial point of optimization lies long before the data processing during the planning of the capture stage itself. For the construction of the stage, we have found that glossy materials should be avoided wherever possible. For example, carpet instead of hardwood flooring creates a diffuse surface and has the added benefit of improved safety against slipping. Diffuse materials should also be preferred for the housing of cameras and scaffolding holding the cameras and lights. It is also beneficial to block the view from the capture stage into the room. This could be done with fabric or matted plexiglass. The goal is a mostly uniform background color. When choosing lights, it is important to consider that small details like hair will disappear when placed in front of bright lights. Spreading the lighting out over a larger area or customizing the light positions according to the capture subject may lead to improvements. Some issues, however, like soft shadows, cannot be solved through the setup and need to be addressed through software.
@@ 4.3 Student-Teacher Learning for Fast Matting
A student-teacher framework is ideal for addressing the performance gap between offline and real-time processing. As shown in the second phase in Fig. 4, the high-capacity network, that has been fine-tuned on the coarse annotations (Sec. 4.2), serves as the teacher model, generating high-quality alpha masks on a larger dataset of unlabeled capture-stage data. These teacher-generated outputs are then used as training data to fine-tune a lightweight, real-time student network. This allows the student model to replicate the structural details that the fine-tuned teacher model has captured while preserving the computational efficiency required for real-time applications. We observed that directly training the lightweight student model on scribble-labeled data reduces its ability to capture fine details, as shown in Sec. 5.2.
@@ 4.4 Quality Control
We found that while the previous approach works, it can still be somewhat error-prone as it lacks proper and quantitative quality control. Therefore, we propose to introduce a validation mechanism using a supervisor model to evaluate both the offline and the real-time methods. The requirements of this model are slightly different, as quality becomes the highest priority, while computational cost and effort required by the user are less important. Given those requirements it makes sense to use a trimap-based matting approach, effectively limiting the range of errors within a small window around the edges, which can typically be near-perfectly retrieved by those methods, at the cost of much longer manual labeling and processing times.
5 Evaluation
@@
Hardware Setup
Our capture stage features $40$ $24.5$ megapixel RGB cameras and $46$ video lights affixed to cylindrical scaffolding with a diameter of $5.25$ meters. The cameras are capable of capturing $75$ frames per second with a sensor depth of $12$ bits. Subjects in the stage are lit by all video lights and captured by all cameras simultaneously. The brightness of the lights in the stage can be controlled, and the camera positions are static, therefore the backgrounds of the cameras do not change and can be pre-captured.
@@
For evaluation, we modified ViTMatte-S [28], referred to as BgViTMatte, which serves as the offline teacher model in our experiments. This included the substitution of the input trimap images with background images, introducing Gaussian noise with a standard deviation of up to $0.1$ during training to adapt the model to our particular problem setup, and training it on Adobe Deep Image Matting [27]. The impact of the noise augmentation is demonstrated in Fig. 5.
@@
We trained BgViTMatte for $2000$ iterations with a batch size of $16$, using an initial learning rate of $5\cdot 10^{-5}$, scheduled to $2.5\cdot 10^{-5}$ after $600$ iterations.
The refined teacher model generates ground truth masks for a new capture stage dataset comprising $20$ scenes with $29$ different camera-views each ($580$ images in total), that are used to fine-tune the student model, BackgroundMattingV2 (BGMV2) [14] with a MobileNetV2 backbone, for real-time performance. BGMV2 employs a two-stage architecture: a base model that generates a coarse mask from a fore- and background image, and a second model that improves fine details along the mask’s borders. We fine-tuned the base model for $5$ epochs, and the combined model for an additional $10$ epochs using the capture stage dataset and the generated ground truth masks. We have kept all other hyperparameters and losses unchanged, and refer to [14] for more details. Validation is conducted using DiffMatte [5] to create a high-detailed benchmark of 14 high-resolution masks based on hand-drawn trimaps from various scenes and camera positions in the capture stage.
@@
Method MSE↓ SAD↓ Grad↓
$\cdot 10^{-4}$ $\cdot 10^{-3}$ $\cdot 10^{-5}$
BgViTMatte $7.805$ $5.760$ $4.221$
BgViTMatte* $\mathbf{4.978}$ $\mathbf{1.415}$ $3.978$
BGMV2 [14] $20.694$ $3.564$ $4.772$
BGMV2* [14] $8.054$ $2.032$ $\mathbf{3.783}$
Table 1: Comparison of teacher (BgViTMatte) and student model (BGMV2), with * indicating task-specific refinements.
@@
5.1 Quantitative and Qualitative Evaluation
To evaluate our approach in comparison to the baseline, we measured the performance of the teacher model on the hold-out validation set, in both its fine-tuned and base configurations, and the student model, trained on data generated by the fine-tuned BgViTMatte in Tab. 1. The table demonstrates the improved performance in terms of the sum of absolute difference (SAD), mean squared error (MSE), and spatial-gradient metric (Grad) of the refined teacher model compared to its base configuration, as well as similar improvements for the student model trained on the teacher-generated data.
To further analyse the impact of the task-specific refinements, we performed a qualitative comparison of both the teacher and student models before and after these refinements. From Fig. 6 we can observe that the teacher model tends to produce false positives in its base configuration, which is mitigated through the refinement. Meanwhile, the student model rather struggles to capture all foreground regions accurately in its base setup. Fine-tuning on the teacher-generated data, however, significantly improves its performance by more precisely outlining foreground elements and reducing errors.
Figure 7: Qualitative comparison of novel views from InstantNGP using the original alpha masks generated by BGMV2 [14] (left) and our improved masks (right). The erroneous matting of the baseline not only cuts out part of the content but also results in an unfaithful color representation to compensate for the wrong inputs.
5.2 Scribble-Based Training of the Student
Fine-tuning the real-time student model BGMV2 [14] directly on sparse scribble annotations is challenging because its second stage (mask-refinement network) is inherently unsuited for training on sparse annotations. To work around this, we first fine-tuned only the first stage of BGMV2 on the hybrid dataset combining our scribble annotations with the Adobe Image Matting Dataset, while keeping the refinement network frozen. We observed an MSE of $89.2\cdot 10^{-4}$ and SAD of $10.4\cdot 10^{-3}$. Next, we separately fine-tuned the mask-refinement network on the Adobe Image Matting Dataset, which reduced the error to an MSE of $27.8\cdot 10^{-4}$ and an SAD of $57.5\cdot 10^{-4}$. Despite this improvement, training the refinement network separately does not fully exploit the potential of our student-teacher approach.
Finally, by integrating the full student-teacher framework we enable the student (real-time) model to learn from the teacher’s high-resolution alpha masks. This joint training yields outputs that retain real-time performance while matching the teacher’s mask quality.
5.3 Learning Capacity Student Model
To assess the student model’s learning capability, we additionally compared its predicted masks to those generated by the teacher model on a validation dataset of $519$ images taken in the capture stage. Specifically, we evaluated the baseline student model BGMV2 [14] with its fine-tuned counterpart. Tab. 2 demonstrate the improved alignment between the student and teacher models after fine-tuning.
MSE↓ SAD↓ Grad↓

### v2

# Capture Stage Matting: Challenges, Approaches, and Solutions for Offline and Real-Time Processing

@@ Abstract
Capture stages are high-end sources of state-of-the-art recordings for downstream applications in movies, games, and other media. One crucial step in almost all pipelines is matting, i.e., separating captured performances from the background. While common matting algorithms deliver remarkable performance in other applications like teleconferencing and mobile entertainment, we found that they struggle significantly with the peculiarities of capture stage content. The goal of our work is to share insights into those challenges as a curated list of these characteristics along with a constructive discussion for proactive intervention and present a guideline to practitioners for an improved workflow to mitigate unresolved challenges. To this end, we also demonstrate an efficient pipeline to adapt state-of-the-art approaches to such custom setups without the need for extensive annotations, both offline and real-time. For an objective evaluation, we introduce a validation methodology using a state-of-the-art diffusion model to demonstrate the benefits of our approach.
@@ 1 Introduction
Matting is a fundamental technique in the field of computer vision, playing a key role in tasks requiring precise foreground-background separation. Mathematically, matting can be expressed by a linear combination of fore- and background images $F$ and $B$, enabling an alpha blending between foreground and background,
@@
Over the last decade, deep learning methods have driven the development of state-of-the-art matting techniques capable of fine-grained transparency estimation, enabling the application of matting techniques in a wide range of fields, including augmented reality and film production. An important application domain lies within capture stages that provide multi-camera setups for recording high-quality data in controlled environments. Robustness of matting is essential for reconstructing high-fidelity content, as large volumes of multi-camera data make per-frame supervision infeasible. However, achieving such matting in these setups remains a significant challenge, as state-of-the-art methods [5, 30, 25] typically rely on predefined trimaps or large-scale labeled datasets, both impractical for capture stages: Trimaps demand highly accurate per-frame generation, and existing datasets generalize poorly to these environments, often leading to errors.
Fortunately, capture stages characteristically feature a static background and controlled lighting, which enable the capture of the empty stage as a reference for matting, providing additional information to support the process. However, this approach introduces new challenges, as reflective materials (e.g., structures within the stage) and soft shadows cast by the subject can significantly alter the appearance of the stage in each shot, complicating the matting workflow, especially given the persistent lack of high-quality training data.
In this work, we address these challenges by discussing the potential failure cases in capture stage setups, providing a guideline for improved matting in these environments, and improving deep matting techniques through a pipeline specifically designed for such capture stage setups. In particular, we place a special focus on real-time matting, which is highly relevant for streaming and telepresence applications. In this context, we discuss a two-phase pipeline based on a student-teacher concept. We modified an offline matting model to apply background matting and adapted it to process capture stage data using coarse user-drawn annotations in the first phase. This adapted model serves as a teacher, whose refined outputs guide a real-time matting model to generate high-quality mattes for capture stage data by learning from the refined outputs of the teacher. We evaluate both models using a hand-labeled trimap-based diffusion model on a hold-out validation set. In summary, our contributions are:
@@
  • We validate our approach against a precise, trimap-based matting method and present the effects of accurate matting in a NeRF-based downstream application, showing improvements in rendering quality (Sec. 5)
@@ 2 Related Work
With the rise of deep learning, Xu et al. [29] introduced a deep image matting model and released a dataset enabling training of models for alpha mask prediction. Since then, numerous works applied learning methods for deep image matting, many of which rely on pre-defined trimaps [30, 5], that often present a bottleneck due to the manual labeling required for their creation. Zhang et al. [33] addressed this challenge by learning trimaps from sparse user input, while others [17] developed a student-teacher model, letting a trimap-free student model learn from a trimap-guided teacher. Hachmann et al. [4] replaced the trimap completely by predicting the matting based on successive images of a scene with changing background colors. To facilitate the matting task, some works built upon segmentation models [11, 31], or used them for additional annotation [8], while another line of research used attention mechanisms and transformer models for matting [30, 2], e.g. by using cross-attention to capture object contours [28]. As diffusion models gained popularity, they have been increasingly utilized in the field of deep image matting [27, 3].
@@
Several works proposed real-time approaches for image and video matting to support certain downstream applications like live-streaming. In this context, Lin et al[14] presented a real-time matting model that uses additional background information for image matting, while other methods [7, 21, 15] solely rely on the image containing the foreground object. Among these methods, Ke et al[7] proposed ModNet, a real-time matting model specifically designed for portrait matting, addressing flickering issues caused by the lack of temporal coherence in video matting. BiMatting [21] applies mask binarization to meet real-time requirements, while a separate recurrent-network solution [15] enhances robustness for continuous video matting. In this context, Huynh et al[6] proposed a mask-guided method for human instance matting, while Lin et al[13] developed a transformer-based approach for matting humans in unstructured video environments. For a comprehensive overview of deep image matting methods, please refer to the survey by Li et al[10].
@@ 3 Problem Setup and Challenges
$I_{i}=\alpha_{i}F_{i}+(1-\alpha_{i})\,\bigl(B_{i}+\delta_{B}(F)\bigr)+\varepsilon_{i},\quad\alpha_{i}\in[0,1],$ (2)
whereby $\delta_{B}(F)$ represents alterations to $I_{i}$ caused by the influence of the foreground subject’s shadows, reflections, or color bleed, while $\varepsilon$ represents sensor noise and other unmodeled artifacts.
Even in highly controlled studio conditions, minor deviations between the “ideal” background $B$ and the actual recordings of the stage with a foreground subject present are inevitable. These background alterations in $I$ are reinforced in particular by reflective materials such as the housings of cameras, lenses or metal frames, which reflect light and the subject. Shadows cast by the subject further alter the appearance of the background, whereby soft shadows occur especially at the feet of a subject, as shown in Fig. 2. Reflections and shadows make the pre-captured background information unreliable, leading to ambiguity in distinguishing the fore- and background. These deviations in the background can substantially affect the accuracy of alpha estimation, as the matting algorithm may fail to attribute background changes correctly to either the foreground layer $F$ or the background layer $B+\delta_{B}(F)$.
Like all camera setups, capture stages are subject to noise $\varepsilon$, which effectively alters the appearance of the background in each recording. Noise degrades the edges of objects, making pixel-accurate distinctions difficult, misleading a matting network into making incorrect predictions. While longer exposure times reduce noise, they also increase undesirable motion blur, particularly when capturing a moving person. In a capture stage setup, this can either require uncomfortably bright lights or high gain levels which further increase the noise and create more pronounced differences in the background.
@@
Figure 3: Ambiguities in foreground and background information can be caused by lighting in the background (left) and dark objects against similarly dark backgrounds (right).
@@ 4.1 Start with a Better Setup
Perhaps the most crucial point of optimization lies long before the data processing during the planning of the capture stage itself. For the construction of the stage, we have found that glossy materials should be avoided wherever possible. For example, carpet instead of hardwood flooring creates a diffuse surface and has the added benefit of improved safety against slipping. Diffuse materials should also be preferred for the housing of cameras and scaffolding holding the cameras and lights. It is also beneficial to block the view from the capture stage into the room. This could be done with fabric or frosted plexiglass. The goal is a mostly uniform background color. When choosing lights, it is important to consider that small details like hair will disappear when placed in front of bright lights. Spreading the lighting out over a larger area or customizing the light positions according to the capture subject may lead to improvements. Some issues, however, like soft shadows, cannot be solved through the setup and need to be addressed through software.
@@ 4.3 Student-Teacher Learning for Fast Matting
A student-teacher framework is ideal for addressing the performance gap between offline and real-time processing. As shown in the second phase in Fig. 4, the high-capacity network that has been fine-tuned on the coarse annotations (Sec. 4.2), serves as the teacher model, generating high-quality alpha masks on a larger dataset of unlabeled capture-stage data. These teacher-generated outputs are then used as training data to fine-tune a lightweight, real-time student network. This allows the student model to replicate the structural details that the fine-tuned teacher model has captured while preserving the computational efficiency required for real-time applications. We observed that directly training the lightweight student model on scribble-labeled data reduces its ability to capture fine details, as shown in Sec. 5.3, which makes intuitive sense, since scribbles usually avoid boundaries, depriving the model of the fine-grained edge information necessary for accurate boundary reconstruction, highlighting the need for a student-teacher approach.
@@ 4.4 Quality Control
Although the previous approach already improves robustness, it remains somewhat error-prone due to the absence of quantitative quality control. To address this limitation, we propose to introduce a validation mechanism using a supervisor model to evaluate both the offline and the real-time methods. The requirements of this model are slightly different, as quality becomes the highest priority, while computational cost and effort required by the user are less important. Given those requirements it makes sense to use a trimap-based matting approach, effectively limiting the range of errors within a small window around the edges, which can typically be near-perfectly retrieved by those methods, at the cost of much longer manual labeling and processing times.
5 Experimental Evaluation and Analysis
@@
5.1 Experimental Setup
We describe the experimental setup, including hardware configuration, model adaptations, and training procedures.
Hardware Configuration
Our capture stage features $40$ RGB cameras ($24.5$ megapixels each, downsampled $2\times$ for most applications) and $46$ video lights affixed to cylindrical scaffolding with a diameter of $5.25$ meters. The cameras are capable of capturing $75$ frames per second with a sensor depth of $12$ bits. Subjects in the stage are lit by all video lights and captured by all cameras simultaneously. The brightness of the lights in the stage can be controlled, and the camera positions are static, therefore the backgrounds of the cameras do not change and can be pre-captured.
@@ 5.1 Experimental Setup
For evaluation, we modified ViTMatte-S [30], referred to as Bg-ViTMatte, which serves as the offline teacher model in our experiments. This included the substitution of the input trimap images with background images, introducing Gaussian noise with a standard deviation of up to $0.1$ during training to adapt the model to our particular problem setup, and training it on the Adobe Deep Image Matting dataset [29]. The positive impact of the noise augmentation is demonstrated in Fig. 5.
@@
We trained Bg-ViTMatte for $2000$ iterations with a batch size of $16$, using an initial learning rate of $5\cdot 10^{-5}$, scheduled to $2.5\cdot 10^{-5}$ after $600$ iterations. The refined teacher model generates pseudo-labels for a new capture stage dataset comprising $20$ scenes with $29$ different camera-views each ($580$ images in total), that are used to fine-tune the student model, BackgroundMattingV2 (BGMV2) [14] with a MobileNetV2 backbone, for real-time performance. BGMV2 employs a two-stage architecture: a base model that generates a coarse mask from a fore- and background image, and a second model that improves fine details along the mask’s borders. We fine-tuned the base model for $5$ epochs, and the combined model for an additional $10$ epochs using the capture stage dataset and the generated pseudo-labels. We have kept all other hyperparameters and losses unchanged, and refer to [14] for more details. Validation is conducted using DiffMatte [5] to create a high-detailed benchmark of 14 high-resolution masks based on hand-drawn trimaps from various scenes and camera positions in the capture stage.
@@
Figure 6: The ratio of sparsely annotated data during training the teacher model significantly affects performance. Training exclusively on scribbled data is not included in the plotted range, resulting in an MSE of $78\cdot 10^{-4}$ and an SAD of $11\cdot 10^{-3}$.
@@
5.2 Baseline Methods and Comparisons
To evaluate the performance of our proposed approach, we compare against a set of baseline methods quantitively, spanning both real-time and high-quality matting models. These include recent architectures that operate either with or without background input.
Quantitative Results
We evaluate the performance of baseline matting approaches on the hold-out validation set using three standard metrics: mean squared error (MSE), sum of absolute differences (SAD), and the spatial-gradient error (Grad), all computed over full-resolution images. The results are summarized in Tab. 1.
First, we establish lower-bound baselines with real-time matting methods that operate without background input, relying solely on RGB data, in the top section of Tab. 1 These models illustrate the challenge of capture-stage matting when explicit scene context is unavailable. As expected, the absence of background cues results in consistently lower accuracy, confirming that for capture-stage applications, matting without auxiliary input is inherently limited – particularly under real-time constraints.
The middle section presents matting methods that operate with additonal background input. Several of these are adapted variants of existing models (prefixed with Bg-) originally designed for RGB or RGB+trimap inputs but modified to accept an additional background image alongside the RGB input. This additional cue consistently improves performance; for example, augmenting DAMNet [26] with background input reduces MSE by more than an order of magnitude (a factor of $14.6$). Since capture-stage data differs from standard open datasets due to lighting, reflections, and shadows, it is essential to incorporate such data into the learning process. In this setting, adding domain adaptation (da) to Bg-DAMNet yields only a modest MSE reduction of $18\%$. In contrast, our scribble-based student–teacher strategy adapts naturally to capture-stage conditions by exploiting additional non-composited training images, resulting in a substantial 61% MSE improvement. This aligns with prior findings for in-the-wild matting, where improved training data was shown to significantly enhance performance [9].
Finally, the bottom section reports results for our proposed models. Bg-ViTMatte, fine-tuned on a hybrid dataset using sparse labels, is exposed to capture-stage data during training. This model achieves the best overall MSE and SAD scores, reducing MSE by $36.2\%$ compared to its non-hybrid counterpart, highlighting the benefits of a hybrid-data training. The distilled BGMV2* model attains the lowest Grad error while maintaining competitive accuracy at $152$ FPS on an NVIDIA GeForce RTX 4070 Ti SUPER. This confirms that our approach scales seamlessly from offline to real-time settings, while avoiding the heavy annotation requirements of conventional training pipelines.
Method MSE ($\downarrow$) SAD ($\downarrow$) Grad ($\downarrow$)